OpenStudy (anonymous):

Solve 14-13x=-3x^2 by using the quadratic formula.

OpenStudy (bahrom7893):

didnt i answer this?

OpenStudy (anonymous):

...it was kinda confusing at the last part.

OpenStudy (anonymous):

3x^2-13x+14=0 3x^2-6x-7x+14=0 3x(x-2)-7(x-2)=0 (3x-7)(x-2)=0 Equate each part in brackets to 0 so, 3x-7=0 3x=7 x=7/3 x-2=0 x=2 Solutions are 2 or 7/3